OpenStudy (anonymous):

Given the coordinates of the vertices of triangle RST are R(-3, 5), S(4, 5) , and T(4, -2). What is the measure of side RT to the nearest tenth? 7.0 9.9 8.2 6.8

OpenStudy (nikato):

do you know the distance formula?

OpenStudy (anonymous):

d/t?

OpenStudy (nikato):

\[\sqrt{(x _{1}-x _{2})^{2}+(y _{1}-y _{2})^{2}}\]

OpenStudy (anonymous):

Oh yes! I remembered.

OpenStudy (nikato):

now plug int he points for R and T

OpenStudy (anonymous):

On it!

OpenStudy (anonymous):

42

OpenStudy (nikato):

how did you get 42?

OpenStudy (nikato):

\[\sqrt{(-3-4)^{2}+[5-(-2)]^{2}}\]

OpenStudy (anonymous):

That's what I did :)

OpenStudy (nikato):

\[=\sqrt{(-7)^{2}+7^{2}}\]

OpenStudy (nikato):

\[=\sqrt{49+49}\]

OpenStudy (anonymous):

9.8?

OpenStudy (nikato):

you shouldve gotten 9.899... and round up to 9.9

OpenStudy (anonymous):

@nikato do you have another time for a problem?

OpenStudy (nikato):

yea. sure

OpenStudy (anonymous):

I did this. But not quite sure which one is the answer. I think it was A.

OpenStudy (anonymous):

Or C..

OpenStudy (nikato):

no. what ratio did you get?

OpenStudy (anonymous):

I got 3/22 or 3:22

OpenStudy (anonymous):

As I remembered. You want me to find the question?

OpenStudy (anonymous):

I'll tag you in the link of it.

OpenStudy (nikato):

yes. so its A. because they have the ratio. PQ and NM are paraellel

OpenStudy (nikato):

and find what question?

OpenStudy (anonymous):

Oh just to show you the same question I asked before. Thank you again!

OpenStudy (nikato):

no problem